OpenStudy (anonymous):

help me please. thanks Given f(x) = x + 2 and g(x) = x^2 – 4, find the function (g/f)(x).

OpenStudy (anonymous):

algebra 2?

OpenStudy (anonymous):

yes

OpenStudy (dtan5457):

Never done a divison composite function...shouldn't be hard though..give me a few minutes.

OpenStudy (dtan5457):

I'm pretty sure what that means is to put f(x) into g(x) (replace x values, then divide the result as g(x)/f(x)

OpenStudy (dtan5457):

\[g(x)=x^2-4\] \[f(x)=x+2\] \[(x+2)^2-4=x^2+4x\]

OpenStudy (dtan5457):

that should of (g o f)x

OpenStudy (dtan5457):

\[\frac{ x^2+4x }{ x+2 }=\frac{ x(x+4) }{ x+2 }\]

OpenStudy (dtan5457):

that i believe is (g/f)(x)..

OpenStudy (dtan5457):

@Sadworld

jimthompson5910 (jim_thompson5910):

\[\Large (g/f)(x) = \frac{g(x)}{f(x)}\] \[\Large (g/f)(x) = \frac{x^2-4}{x+2}\] \[\Large (g/f)(x) = \frac{x^2-2^2}{x+2}\] \[\Large (g/f)(x) = \frac{(x-2)(x+2)}{x+2}\] I'm sure you see what to do from here

OpenStudy (dtan5457):

ooh

OpenStudy (anonymous):

thank you both (: